# The role of the irreducible mass in repetitive Penrose energy extraction processes in a Kerr black hole

### v1

@@ Abstract
The concept of the irreducible mass ($M_{\rm irr}$) has led to the mass-energy ($M$) formula of a Kerr-Newman black hole (BH), in turn leading to an expression for its surface area $S=16\pi M_{\rm irr}^{2}$. This allowed as well the coeval identification of the reversible and irreversible transformations. The concepts of extracted and extractable energy soon followed. This new conceptual framework avoids inconsistencies recently evidenced in a repetitive Penrose process. We consider repetitive decays in the ergosphere of an initially extreme Kerr BH and show that these processes are highly irreversible. For each decay, including the BH capture of the particle falling into the horizon, the increase of the irreducible mass is much larger than the extracted energy. Correspondingly, the horizon radius increases. The energy extraction process saturates and stops when the horizon reaches the decay radius. Thus, reaching a final non-rotating Schwarzschild BH state by this accretion process is impossible. We have assessed such processes for three selected decay radii, $r=1.2M$, $1.5M$, and $1.9M$. The energy extracted until saturation is only $2.75\%$, $4.54\%$, and $2.45\%$ of the BH initial mass, while the extractable energy is reduced by $25\%$, $56\%$, and $92\%$. The BH rotational energy is mostly converted into irreducible mass, which increases to $75\%$, $83\%$, and $95\%$ of the initial BH mass. Therefore, evaluating the effect of the irreducible mass increase in the energy extraction processes in the Kerr BH ergosphere is now mandatory.
@@ I Introduction
In a recent article [1] (hereafter Paper I), we have revisited the equations describing a single Penrose process of a test particle $\mu_{0}$ that decays into two particles $\mu_{1}$ and $\mu_{2}$ in the ergosphere of an extreme Kerr black hole (BH) with mass $M$ and spin $\hat{a}=a/M=1$. We assumed $\mu_{0}=10^{-2}M$, fulfilling the condition of being a test particle. We addressed the decay process in three radii inside the BH ergosphere: $r=1.2M$, $1.5M$, and $1.9M$. The basic assumption to describe the process was that the decay of the test particles would occur at the turning points of their trajectories. We have further shown that to identify a specific process, we could adopt two additional conditions: one that the angular momentum of the particle $\mu_{1}$ entering the BH could be assumed in all three decay processes to be equal to $p_{\phi 1}=-19.434\,\mu_{1}M$ and that the ratio of the two particles $\mu_{2}/\mu_{1}=0.78345$. This choice has the advantage, shown in Paper I, of leading to a working solution of the decay equations (see Section II therein). Still, in Paper I, we also showed that if one turns from a single Penrose process, which is linear in the BH mass and angular momentum, an attempt to apply this linear process into a sequence of repetitive events leads to a violation of energy conservation. Thus, a nonlinear treatment of the mass and angular momentum becomes necessary. The topic of this new paper is considering the nonlinear relation between the BH mass, angular momentum, and the irreducible mass [2]. The latter separates the BH transformations into reversible ones that keep constant the irreducible mass and irreversible ones in which it monotonically increases.
@@
To approach this problem in the most general formulation, we express in Section II the general equations for the mass-energy of a Kerr Newman metric characterized by mass, charge, angular momentum, and the irreducible mass [3, 4]. We also give the upper limits to the charge and spin parameters and the expression of the horizon surface area. In addition to these classic results, we also express the inverse relation relating the irreducible mass to the BH mass, spin, and charge. In total generality, fulfilling the above limitations, we also introduce the extracted energies presented in Paper I and the novel concept of extractable energy. We recall how the maximum extractable energy cannot exceed $29.3\%$ of the BH mass of an extreme Kerr BH and $50\%$ of the mass of an extremely charged Reissner-Nordström BH. Intermediate values apply for other BH parameters of the Kerr-Newman BH.
@@
In Section III, we extend the analytic solution of the Penrose Process, introduced in Paper I for the extreme BH, to any spin value. We show the appearance of the irreducible mass role in the first step, starting from an extreme Kerr BH, again for the same fixed assumptions (see Table 1). The fundamental result is that a Penrose process is far from being linear, and the correct theoretical approach shows it has a highly nonlinear and irreversible behavior. Even in this first example, the increase of the irreducible mass is much larger than the extracted energy.
@@
Section IV shows a repetitive nonlinear process where we formulate the iterative process going from $n-1$ to $n$. Table 2 and Fig. 1 summarize the explicit example of decays at $r=1.2M$. We finally proceed to the conclusions in Section V.
@@
It is clear that due to the contribution of the irreducible mass strongly increasing at each step, the energy extraction is saturated after a finite number of steps in each case. The final value of the extractable energy remains very close to the initial value, or it is reduced owing to the conversion of rotational energy into irreducible mass. In the former case, the remaining extractable energy can be extracted by a reversible electrodynamical process, as shown in the third article in this series (Rueda and Ruffini, to be submitted).
@@ II The irreducible mass
where we have define the dimensionless spin $\hat{a}\equiv L/M^{2}$ and charge $\hat{e}\equiv e/M$.
@@
which monotonically increases for irreversible transformations and is constant otherwise. For an infinitesimal change of the irreducible mass $\delta M_{\rm irr}$, the surface area increases by
$\delta S=32\pi M_{\rm irr}\delta M_{\rm irr}\ .$ (5)
@@
With the irreducible mass, we introduce the concept of BH extractable energy
@@ III The first appearance of the irreducible mass increase in the Penrose process
We now specialize in the Kerr BH, i.e., $e=0$. In this case, the system of equations of the Penrose process is given by the energy and momentum conservation
$\mu_{0}\hat{E}_{0}$ $=\mu_{1}\hat{E}_{1}+\mu_{2}\hat{E}_{2},$ (8a)
$\mu_{0}\hat{p}_{\phi 0}$ $=\mu_{1}\hat{p}_{\phi 1}+\mu_{2}\hat{p}_{\phi 2},$ (8b)
$\mu_{0}\hat{p}_{r0}$ $=\mu_{1}\hat{p}_{r1}+\mu_{2}\hat{p}_{r2},$ (8c)
and the request that the decay point be a triple turning point, i.e., the vanishing of the radial momentum of the three particles therein
$\hat{p}_{r0}$ $=-\left[(\hat{E}_{0}^{2}-1)\hat{r}^{3}+2\hat{r}^{2}\right.$
$+\left.(\hat{a}^{2}\hat{E}_{0}^{2}-\hat{p}_{\phi 0}^{2}-\hat{a}^{2})\hat{r}+2(\hat{E}_{0}\hat{a}-\hat{p}_{\phi 0})^{2}\right]^{1/2},$ (9a)
$\hat{p}_{r1}$ $=-\left[(\hat{E}_{1}^{2}-1)\hat{r}^{3}+2\hat{r}^{2}\right.$
$+\left.(\hat{a}^{2}\hat{E}_{1}^{2}-\hat{p}_{\phi 1}^{2}-\hat{a}^{2})\hat{r}+2(\hat{E}_{1}\hat{a}-\hat{p}_{\phi 1})^{2}\right]^{1/2}=0,$ (9b)
$\hat{p}_{r2}$ $=\left[(\hat{E}_{2}^{2}-1)\hat{r}^{3}+2\hat{r}^{2}\right.$
$+\left.(\hat{a}^{2}\hat{E}_{2}^{2}-\hat{p}_{\phi 2}^{2}-\hat{a}^{2})\hat{r}+2(\hat{E}_{2}\hat{a}-\hat{p}_{\phi 2})^{2}\right]^{1/2}=0.$ (9c)
Assuming the incoming particle starts at rest from infinity ($\hat{E}_{0}=1$), that the decay occurs at some given $\hat{r}$, and for given $\hat{p}_{\phi 1}$ and the outgoing to infalling particle mass-ratio $\mu_{2}/\mu_{1}$, the system of equations can be solved analytically for the other variables. The solution is
$\hat{p}_{\phi 0}$ $=\frac{2\hat{a}-\sqrt{2\hat{r}\Delta}}{2-\hat{r}},$ (10a)
$\hat{E}_{1}$ $=\frac{2\,\hat{a}\,\hat{p}_{\phi 1}+\sqrt{\rho+\hat{r}^{2}\Delta\,\hat{p}_{\phi 1}^{2}}}{R},$ (10b)
$\hat{E}_{2}$ $=\frac{1}{\tilde{\mu}_{2}}-\frac{\mu_{1}}{\mu_{2}}\hat{E}_{1},$ (10c)
$\hat{p}_{\phi 2}$ $=\frac{1}{\tilde{\mu}_{2}}\hat{p}_{\phi 0}-\frac{\mu_{1}}{\mu_{2}}\hat{p}_{\phi 1},$ (10d)
$\tilde{\mu}_{1}$ $=\frac{\rho}{K+\sqrt{K^{2}-\rho^{2}[1-(\mu_{2}/\mu_{1})^{2}]}},$ (10e)
where $\tilde{\mu}_{1}=\mu_{1}/\mu_{0}$, $\tilde{\mu}_{2}=\mu_{2}/\mu_{0}$, and
$\Delta$ $=\hat{r}^{2}-2\hat{r}+\hat{a}^{2},\quad R=\hat{r}\Delta+2(\hat{r}^{2}+\hat{a}^{2}),\quad\rho=\hat{r}\,\Delta\,R$ (11a)
$K$ $=\sqrt{(\rho+\hat{r}^{2}\Delta\hat{p}_{\phi 0}^{2})(\rho+\hat{r}^{2}\Delta\hat{p}_{\phi 1}^{2})}-\hat{r}^{2}\Delta\hat{p}_{\phi 0}\hat{p}_{\phi 1}.$ (11b)
@@
$\Delta E_{{\rm extractable},0}=\Delta M_{0}-\Delta M_{{\rm irr}_{0}},$ (14)
@@
Assuming an incoming particle that decays at $\hat{r}=1.2$, an angular momentum of the infalling particle $\hat{p}_{\phi 1}=-19.434$, and the outgoing to infalling particle mass-ratio $\mu_{2}/\mu_{1}=0.78345$, we obtain from Eqs. (10) and (11) the values of the other system parameters for a given value of the decay point. Table 1 summarizes the results for decays at $\hat{r}=1.2$, $1.5$, and $1.9$, assuming $\mu_{0}=10^{-2}M_{0}$.
For instance, in the case $\hat{r}=1.2$, the extracted energy, $E_{{\rm extracted},0}=|\Delta M_{0}|=0.00145M_{0}$, is much lower than the increase in the irreducible mass, $\Delta M_{{\rm irr},0}=0.0158M_{0}$, which has got a boost of more than ten times larger. This result demonstrates the inefficiency of the process, which converts most BH rotational energy into irreducible mass. Therefore, the process is highly irreversible. We shall return to this point below in a definite way.
Table 1: Summary of Penrose process parameters occurring inside the ergosphere of extreme Kerr BH of mass $M_{0}$ and angular momentum $L_{0}=M_{0}^{2}$. The BH captures the decay-product particle with negative energy $E_{1}=\hat{E}_{1}\mu_{1}$ and negative angular momentum $p_{\phi 1}=\hat{p}_{\phi 1}\mu_{1}M_{0}$. We list the results for the decay occurring at three selected positions: $\hat{r}=1.2$, $1.5$, and $1.9$. We choose $\hat{E}_{0}=1$, $\hat{p}_{\phi 1}=-19.434$, $\mu_{2}/\mu_{1}=0.78345$ and obtain the other parameters from the conservation equations presented in Paper I and in Section III. We adopt $\mu_{0}=10^{-2}M_{0}$ for all cases.
$\hat{r}$ $\frac{\mu_{1}}{\mu_{0}}$ $\frac{\mu_{2}}{\mu_{0}}$ $\hat{E}_{1}$ $\hat{E}_{2}$ $\hat{p}_{\phi 2}$ $\frac{M}{M_{0}}$ $\hat{a}$ $\frac{\Delta M_{{\rm irr},0}}{M_{0}}$ $\frac{E_{{\rm extracted},0}}{M_{0}}$ $\frac{E_{\rm extractable}}{M_{0}}$
$1.2$ $0.0209$ $0.0164$ $-6.936$ $69.642$ $153.872$ $0.9986$ $0.999$ $0.0158$ $0.00145$ $0.276$
$1.5$ $0.0218$ $0.0171$ $-3.520$ $62.995$ $157.485$ $0.9992$ $0.997$ $0.0245$ $0.000768$ $0.267$
$1.9$ $0.0247$ $0.0194$ $-0.501$ $52.326$ $151.737$ $0.9999$ $0.995$ $0.0328$ $0.000124$ $0.260$

### v3

@@ Abstract
The concept of the irreducible mass ($M_{\rm irr}$) has led to the mass-energy ($M$) formula of a Kerr-Newman black hole (BH), in turn leading to an expression for its surface area $S=16\pi M_{\rm irr}^{2}$. This also allowed the coeval identification of the reversible and irreversible transformations. The concepts of extracted and extractable energy soon followed. This new conceptual framework avoids inconsistencies recently evidenced in a repetitive Penrose process. We consider repetitive decays in the ergosphere of an initially extreme Kerr BH and show that these processes are highly irreversible. For each decay, including the BH capture of the particle falling into the horizon, the increase of the irreducible mass is much larger than the extracted energy. Correspondingly, the horizon radius increases. The energy extraction process saturates and stops when the horizon reaches the decay radius. Thus, reaching a final non-rotating Schwarzschild BH state by this accretion process is impossible. We have assessed such processes for three selected decay radii, $r=1.2M$, $1.5M$, and $1.9M$. The energy extracted until saturation is only $2.75\%$, $4.54\%$, and $2.45\%$ of the BH initial mass, while the extractable energy is reduced by $25\%$, $56\%$, and $92\%$. The BH rotational energy is mostly converted into irreducible mass, which increases to $75\%$, $83\%$, and $95\%$ of the initial BH mass. Therefore, evaluating the effect of the irreducible mass increase in the energy extraction processes in the Kerr BH ergosphere is now mandatory.
@@ I Introduction
In a recent article [1] (hereafter Paper I), we have revisited the equations describing a single Penrose process of a test particle $\mu_{0}$ that decays into two particles $\mu_{1}$ and $\mu_{2}$ in the ergosphere of an extreme Kerr black hole (BH) with mass $M$ and spin $\hat{a}=a/M=1$. We assumed $\mu_{0}=10^{-2}M$, fulfilling the condition of being a test particle. We addressed the decay process in three radii inside the BH ergosphere: $r=1.2M$, $1.5M$, and $1.9M$. The basic assumption to describe the process was that the decay of the test particles would occur at the turning points of their trajectories. We have further shown that to identify a specific process, we could adopt two additional conditions: one that the angular momentum of the particle $\mu_{1}$ entering the BH could be assumed in all three decay processes to be equal to $p_{\phi 1}=-19.434\,\mu_{1}M$ and that the ratio of the two particles $\mu_{2}/\mu_{1}=0.78345$. This choice has the advantage, shown in Paper I, of leading to a working solution of the decay equations. Still, in Paper I, it was shown that if one turns from a single Penrose process, which is linear in the BH mass and angular momentum, into a sequence of repetitive events (as has been suggested in some hypothetical applications of the Penrose process, e.g., [2]), keeping the above linearity leads to a violation of energy conservation. Thus, a nonlinear treatment of the mass and angular momentum becomes necessary. The topic of this paper is considering the nonlinear relation between the BH mass, angular momentum, and the irreducible mass [3]. The latter classifies BH transformations into reversible ones that keep the irreducible mass constant and irreversible ones in which it increases.
@@
To approach this problem in the most general formulation, we express first the general equations for the mass-energy of a Kerr-Newman metric characterized by mass, charge, angular momentum, and the irreducible mass [4, 5]. We also give the upper limits to the charge and spin parameters and the expression of the horizon surface area. In addition to these classic results, we express the inverse relation relating the irreducible mass to the BH mass, spin, and charge. In total generality, fulfilling the above limitations, we also introduce the extracted energies presented in Paper I and the novel concept of extractable energy. We recall how the maximum extractable energy cannot exceed $29.3\%$ of the BH mass of an extreme Kerr BH and $50\%$ of the mass of an extremely charged Reissner-Nordström BH. Intermediate values apply for other BH parameters of the Kerr-Newman BH.
@@
Next, we extend the analytic solution of the Penrose Process, introduced in Paper I for the extreme BH, to any spin value. We show the appearance of the irreducible mass role in the first step, starting from an extreme Kerr BH, again for the same fixed assumptions (see Table 1). The fundamental result is that a Penrose process is far from linear, and the correct theoretical approach shows it has a highly nonlinear and irreversible behavior. Even in this first example, the irreducible mass increase is much larger than the extracted energy.
@@
We then show a repetitive nonlinear process where we formulate the iterative process going from $n-1$ to $n$. Table 2 and Fig. 1 summarize the explicit example of decays at $r=1.2M$. We then proceed to the conclusions.
@@
It is clear that due to the contribution of the irreducible mass strongly increasing at each step, the energy extraction is saturated after a finite number of steps in each case. The final value of the extractable energy remains very close to the initial value, or it is reduced owing to the conversion of rotational energy into irreducible mass. In the former case, the remaining extractable energy can be extracted by a reversible electrodynamical process, as shown in a third article (Rueda and Ruffini, submitted).
@@ II The irreducible mass
Throughout, we use geometric units ($c=G=1$) and denote with over hats dimensionless quantities. The dimensionless BH spin and charge are $\hat{a}\equiv L/M^{2}$ and $\hat{e}\equiv e/M$. The dimensionless particle energy is $\hat{E}_{i}=E_{i}/\mu_{i}=\gamma_{i}$ (Lorentz factor), and the radial and angular momentum, $\hat{p}_{ri}=p_{ri}/\mu_{i}$ and $\hat{p}_{\phi i}=p_{\phi i}/(\mu_{i}M)$, respectively.
@@
which for an infinitesimal change $\delta M_{\rm irr}$, increases by
$\delta S=32\pi M_{\rm irr}\delta M_{\rm irr},$ (5)
so it increases for irreversible transformations ($\delta M_{\rm irr}>0$) and constant for reversible ones ($\delta M_{\rm irr}=0$).
@@
With the irreducible mass, we introduce the concept of BH extractable energy [6, 7]
@@ III The first appearance of the irreducible mass increase in the Penrose process
We now specialize in the Kerr BH, i.e., $e=0$, and particle motion on the equatorial plane. In this case, the system of equations of the Penrose process is given by the energy and momentum conservation
$\hat{E}_{0}$ $=\tilde{\mu}_{1}\hat{E}_{1}+\tilde{\mu}_{2}\hat{E}_{2},$ (8a)
$\hat{p}_{\phi 0}$ $=\tilde{\mu}_{1}\hat{p}_{\phi 1}+\tilde{\mu}_{2}\hat{p}_{\phi 2},$ (8b)
$\hat{p}_{r0}$ $=\tilde{\mu}_{1}\hat{p}_{r1}+\tilde{\mu}_{2}\hat{p}_{r2},$ (8c)
where $\tilde{\mu}_{1}=\mu_{1}/\mu_{0}$ and $\tilde{\mu}_{2}=\mu_{2}/\mu_{0}$, and the request that the decay point be a triple turning point of the effective potential of the radial motion [8, 9], i.e., the vanishing of the radial momentum of the three particles therein
$\hat{r}(\hat{E}_{i}^{2}-1)(\hat{r}^{2}+\hat{a}^{2})+2\hat{r}^{2}-\hat{r}\hat{p}_{\phi i}^{2}+2(\hat{E}_{i}\hat{a}-\hat{p}_{\phi i})^{2}=0,$ (9)
where $i=0$, $1$, and $2$. By assuming we know $\hat{E}_{0}$, $\hat{r}$, $\hat{p}_{\phi 1}$, and $\nu\equiv\mu_{2}/\mu_{1}$, the system of equations can be solved analytically for the other variables. The solution is
$\hat{p}_{\phi 0}$ $=\frac{2\hat{a}{\hat{E}_{0}}-\sqrt{\hat{r}\Delta[2{+\hat{r}(\hat{E}_{0}^{2}-1)}]}}{2-\hat{r}},$ (10a)
$\hat{E}_{1}$ $=\frac{2\,\hat{a}\,\hat{p}_{\phi 1}+\sqrt{\hat{r}\Delta(R+\hat{r}\,\hat{p}_{\phi 1}^{2})}}{R},$ (10b)
$\tilde{\mu}_{1}$ $={\frac{\hat{r}\Delta}{K+\sqrt{K^{2}-\hat{r}^{2}\Delta^{2}(1-\nu^{2})}},}$ (10c)
$\hat{E}_{2}$ $={\frac{\hat{E}_{0}}{\tilde{\mu}_{2}}-\frac{\hat{E}_{1}}{\nu}},\quad\hat{p}_{\phi 2}={\frac{\hat{p}_{\phi 0}}{\tilde{\mu}_{2}}-\frac{\hat{p}_{\phi 1}}{\nu}},$ (10d)
where $\tilde{\mu}_{2}=\nu\tilde{\mu}_{1}$, $\Delta=\hat{r}^{2}-2\hat{r}+\hat{a}^{2}$, $R=\hat{r}^{3}+\hat{a}^{2}\hat{r}+2\hat{a}^{2}$, and $K=R\,\hat{E}_{0}\hat{E}_{1}-2\hat{a}(\hat{E}_{0}\hat{p}_{\phi 1}+\hat{E}_{1}\hat{p}_{\phi 0})+(2-\hat{r})\hat{p}_{\phi 0}\,\hat{p}_{\phi 1}$.
@@
$\Delta E_{{\rm extractable},0}=\Delta M_{0}-\Delta M_{{\rm irr},_{0}},$ (13)
@@
Assuming $\hat{p}_{\phi 1}=-19.434$, $\nu=\mu_{2}/\mu_{1}=0.78345$, we obtain from Eqs. (10) the rest of parameters, for given values of $\hat{E}_{0}$ and $\hat{r}$. Table 1 summarizes the results for $\hat{r}=1.2$, $1.5$, and $1.9$, for $\hat{E}_{0}=\gamma_{0}=1$ (at rest at infinity) and $1.5$ (incoming with initial velocity $0.556c$). We use $\mu_{0}=10^{-2}M_{0}$ in all cases.
For instance, in the case $\hat{r}=1.2$ and $\hat{E}_{0}=1$, the extracted energy, $E_{{\rm extracted},0}=|\Delta M_{0}|=0.00145M_{0}$, is much lower than the increase in the irreducible mass, $\Delta M_{{\rm irr},0}=0.0158M_{0}$, which has got a boost of more than ten times larger. Non-zero initial velocities of the incoming particle (see, e.g., the case $\hat{E}_{0}=1.5$) make the situation worse. This result demonstrates the inefficiency of the process, which converts most BH rotational energy into irreducible mass. Therefore, the process is highly irreversible. We shall return to this point below in a definite way.
Table 1: Summary of Penrose process parameters occurring inside the ergosphere of extreme Kerr BH of mass $M_{0}$ and angular momentum $L_{0}=M_{0}^{2}$. The BH captures the decay-product particle with energy $E_{1}=\hat{E}_{1}\mu_{1}<0$ and angular momentum $p_{\phi 1}=\hat{p}_{\phi 1}\mu_{1}M_{0}<0$. We list the results for the decay occurring at three selected positions: $\hat{r}=1.2$, $1.5$, and $1.9$. We choose $\hat{p}_{\phi 1}=-19.434$, $\nu=0.78345$, $\hat{E}_{0}=1.0$ (upper) and $1.5$ (lower) and obtain the other parameters from the conservation equations and the BH parameters from Eqs. (10)–(14). We adopt $\mu_{0}=10^{-2}M_{0}$ for all cases.
$\hat{r}$ $\frac{\mu_{1}}{\mu_{0}}$ $\frac{\mu_{2}}{\mu_{0}}$ $\hat{E}_{1}=\gamma_{1}$ $\hat{E}_{2}=\gamma_{2}$ $\hat{p}_{\phi 2}$ $\frac{M}{M_{0}}$ $\hat{a}$ $\frac{\Delta M_{{\rm irr},0}}{M_{0}}$ $\frac{E_{{\rm extracted},0}}{M_{0}}$ $\frac{E_{\rm extractable}}{M_{0}}$
$\hat{E}_{0}=\gamma_{0}=1$
$1.2$ $0.0209$ $0.0164$ $-6.936$ $69.923$ $153.829$ $0.9986$ $0.9988$ $0.0158$ $0.00145$ $0.276$
$1.5$ $0.0218$ $0.0171$ $-3.521$ $62.995$ $157.483$ $0.9992$ $0.9973$ $0.0250$ $0.000768$ $0.267$
$1.9$ $0.0247$ $0.0194$ $-0.501$ $52.326$ $151.736$ $0.9999$ $0.9954$ $0.0328$ $0.000124$ $0.260$
$\hat{E}_{0}=\gamma_{0}=1.5$
$1.2$ $0.0149$ $0.0117$ $-6.936$ $137.101$ $301.621$ $0.9989$ $0.9992$ $0.0135$ $0.00103$ $0.278$
$1.5$ $0.0151$ $0.0119$ $-3.521$ $131.021$ $327.550$ $0.9995$ $0.9981$ $0.0209$ $0.000533$ $0.271$
$1.9$ $0.0166$ $0.0130$ $-0.501$ $116.086$ $336.646$ $0.9999$ $0.9969$ $0.0270$ $0.000083$ $0.266$